Solve
root of 2x+8=x

Answer by tutorcecilia(2152)

2x+8=x
2x-x+8=x-x [isolate the x]
x+8=0
x+8-8=-8
x=-8
.
check by plugging x=-8 back into the original equation
2x+8=x
2(-8)+8=(-8)
-16-8=-8
-8=-8 [checks out]